Use n=6 subdivision to approximate the value of
∫_0^6▒√(x+2) dx by the trapezoidal rule
1
Expert's answer
2020-10-21T17:24:59-0400

"∫_0^6\sqrt{(x+2)}"

We divide the segment [0;6] into 6 parts and at each node we calculate the value of the integrand

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c:}
 n & 0 & 1&2&3&4&5&6\\ \hline
 x_n & 0 & 1&2&3&4&5&6\\
 \hdashline
 \approx\sqrt{x+2} &1.41&1.73&2&2.24&2.45&2.65&2.83
\end{array}"

The formula for calculating the integral by the trapezoid method

"∫_a^bf(x)dx\approx{h}[\frac{f(x_0)+f(x_n)}{2}+f(x_{1})+...+f(x_{n-1})]\ where \ h =\frac{(b-a)}{n}"

"h=\frac{6-0}{6}=1"

"∫_0^6\sqrt{(x+2)}\approx\frac{1.41+2.83}{2}+1.73+2.0+2.24+2.45+2.65 =13.19"


Answer: "∫_0^6\sqrt{(x+2)}\approx13.19"